# Aligning Translation-Specific Understanding to General Understanding in Large Language Models

### v1

@@ Abstract
Although large language models (LLMs) have shown surprising language understanding and generation capabilities, they have yet to gain a revolutionary advancement in the field of machine translation. One potential cause of the limited performance is the misalignment between the translation-specific understanding and general understanding inside LLMs. To align the translation-specific understanding to the general one, we propose a novel translation process xIoD (Cross-Lingual Interpretation of Difficult words), explicitly incorporating the general understanding on the content incurring inconsistent understanding to guide the translation. Specifically, xIoD performs the cross-lingual interpretation for the difficult-to-translate words and enhances the translation with the generated interpretations. Furthermore, we reframe the external tools of QE to tackle the challenges of xIoD in the detection of difficult words and the generation of helpful interpretations. We conduct experiments on the self-constructed benchmark Challenge-MT, which includes cases in which multiple SOTA translation systems consistently underperform. Experimental results show the effectiveness of our xIoD, which improves up to +3.85 COMET. ††footnotetext: 🖂 means corresponding author.
@@
Yichong Huang, Xiaocheng Feng, Baohang Li, Chengpeng Fu, Wenshuai Huo, Ting Liu, Bing Qin Harbin Institute of Technology Peng Cheng Laboratory {ychuang,xcfeng,baohangli,cpfu,wshuo,tliu,qinb}@ir.hit.edu.cn
@@ 1 Introduction
Recently, with the scaling of model capacities and pre-training data volume, large language models (LLMs) have demonstrated remarkable language understanding and generation, paving the way for a higher level of performance in machine translation Zhao et al. (2023); OpenAI (2023); Jiang et al. (2023); Workshop (2023). However, existing research reports that LLMs have yet to achieve as significant advances in machine translation as they have achieved in other natural language processing fields Hendy et al. (2023); Zhang et al. (2023a); Jiao et al. (2023b); Zhu et al. (2023b).
One potential reason leading to the limited translation performance of LLMs is the misalignment between the general understanding and translation-specific understanding inside LLMs. This claim is based on our discovery that LLMs understand many concepts accurately when explaining directly but often misunderstand these concepts in translation requests, which is illustrated in Fig 1. This paper refers to these concepts as language models’ generalization failures on translation.
@@
In this work, we propose a novel translation process xIoD, which explicitly incorporates the general understanding to guide the translation during inference. Specifically, xIoD first detects the difficult-to-translation words in the source sentence, which could cover the generalization failures intuitively. Next, the LLM is prompted to interpret each difficult word with the target language, i.e., cross-lingual interpretation of difficult words, unleashing the powerful general understanding and aligning this understanding into the target language space. After that, xIoD conducts translation under the guidance of these interpretations. Unlike the CoT-based process mimicking the junior translator to perform word-for-word translation Peng et al. (2023), xIoD is more in accordance with the senior translator owning a higher performance frontier. However, xIoD faces two challenges: (1) LLMs may struggle with the accurate detection of difficult words due to their limitation in self-knowledge Yin et al. (2023), and (2) LLMs are prone to errors and hallucinations in the interpretations Huang et al. (2023a, b), biasing the translation from the original semantics. Therefore, we reframe the external tool of token-level QE Rei et al. (2023) to enhance the detection of difficult words and design a strategy of interpretations quality control to filter hallucinated interpretations based on sentence-level QE Rei et al. (2020).
Towards better identifying the limitations of current MT systems, we proposed the benchmark Challenge-MT via collecting the difficult cases that multiple SOTA systems consistently underperform on multi-year WMT datasets. All experiments are conducted under the few-shot setting, where the demonstrations are constructed automatically in a post-explanation manner. Experimental results on six translation directions (Chinese, Estonian, and Icelandic to/from English) demonstrate the effectiveness of xIoD, which improves up to +3.85 COMET.
@@ 2.1 LLM-based MT
2.2 QE for MT
Quality estimation (QE) for machine translation, i.e., reference-free MT evaluation, aims to predict the quality of the given translation only according to the source sentence, which has shown auspicious correlations with human judgments Rei et al. (2020, 2021). Given a source sentence and the translation , the QE score is denoted as .
@@ 2.2 QE for MT
3 Approach: xIoD
In this section, we first introduce our translation framework xIoD (§3.1). Specifically, xIoD consists of three components: difficult word detection (§3.2), cross-lingual interpretation (§3.3), and interpretation quality control (§3.4). To make the LLM follow the procedure of each component as expected, we adopt the in-context learning strategy and design an automatic method for constructing demonstrations of xIoD (§3.5).
@@ 3.1 Framework
The progress of xIoD is illustrated in Figure 2. Given the source sentence, xIoD first detects the difficult words or phrases in the source sentence. Once the difficult words are identified, xIoD requests the LLM to interpret each difficult word with the target language, unleashing the powerful understanding capability inside the LLM and aligning these understanding into the target language. Finally, to avoid the negative effect of incorrect and useless interpretations, xIoD removes the negative interpretations through the interpretation quality control and outputs the final translation guided by the helpful interpretations. We describe each step in the later.
@@ 3.2 Step-1: Difficult Word Detection
Correct understanding of the difficult words determines largely the quality of the final translation. To make the LLM comprehend these difficult words, it is necessary to recognize them at first. However, finding the translation-specific difficult words remains challenging Lim et al. (2023); Sun (2015). To tackle this challenge, we first conduct a preliminary translation for the given source sentence and then extract the mistranslated words in the source sentence as the difficult words. Concretely, we invent xIoD-I to do this using solely the Intrinsic knowledge of LLMs at first.
xIoD-I.
Given source sentence , xIoD-I first obtains the preliminary translation (also known as draft translation) by prompting the LLM to translate with the in-context learning strategy, which is shown in Eq. (1). Next, the LLM is requested to output the difficult words based on the source sentence and the preliminary translation:
xIoD-I obtains the difficult word list via performing greedy decoding on the LLM:
@@
xIoD-E.
It is frequently observed that the LLM fails to recognize the mistranslated words. Therefore, we devise xIoD-E to boost the detection with the external tool. First, xIoD-E requests the LLM with the same prompt as xIoD-I while performing the temperature sampling for times (). Next, the union of all sampling results is taken as the candidate set of difficult words :
where is the sampling temperature††https://platform.openai.com/docs/api-reference/chat/create#chat-create-temperature, which is set to to capture more candidates.
Finally, xIoD-E annotates each candidate word with its degree of misalignment with respect to the draft translation, which reflects the translation-specific difficulty. To implement this function, we adopt an external tool of token-level QE . As shown in §2.2, token-level QE is originally used to score the mistranslation degree of the given translation span with respect to the source sentence, i.e., where . Differently, we utilize this tool in a dual manner. That is, we use to annotate the misalignment degree of the given source span with respect to the translation, i.e., where . Formally, the misalignment score of each difficult word candidate is calculated as:
Then, xIoD-E selects candidates with misalignment score , where is the hyperparameter named the difficulty threshold. We refer to this procedure as the difficulty-aware selection in Fig 2.
@@ 3.3 Step-2: Cross-Lingual Interpretation
After the difficult words in the source sentence are detected, xIoD lets the LLM generate the interpretation of each difficult word via requesting with the prompt:
@@
Through the generated interpretations, xIoD enhances the translation with LLMs‘ general understanding capability. However, LLMs may generate incorrect or hallucinated interpretations sometimes (e.g., the interpretation of ”崩塌” in Fig 2), which biases the resulting translation from the original semantics. Besides, helpless interpretations that can not provide useful information also pose a risk of disturbing the translation process.
@@ 3.4 Step-3: Interpretation Quality Control
To overcome the potential negative effect of the generated interpretations, xIoD removes the incorrect and useless interpretations through the interpretation quality control (IQC) and outputs the final translation guided by the helpful interpretations.
Concretely, given a set of interpretations , xIoD ablates each interpretation sequentially and uses the remaining interpretations to guide the translation. The interpretation-guided translation is implemented in a fashion of refinement:
@@
where , which is the demonstrations of interpretation-guide translation.
Once the better translation performance is achieved by ablation, which is measured by the QE††We use wmt21-comet-qe-da as the QE scorer. tool due to the unavailable access to the reference translation, the interpretation is removed from and the current translation is taken as the best translation. We detail this process in Alg. 1.
3.5 Demonstrations Synthesis for xIoD
To make the LLM follow the procedure of xIoD as expected, we adopt the ICL strategy. Common practice constructs the demonstrations manually for ICL, necessitating human translators proficient in language pairs for languages. To overcome this considerable cost, we devise a method for synthesizing high-quality demonstrations of xIoD based on the parallel data.
Then, the response is parsed via regular expression to extract the difficult words and interpretations . Next, we remove the noisy interpretations through a process similar to IQC (Alg. 1). The only difference is that the QE metric is replaced with the reference-based COMET Rei et al. (2020) due to the available access to the reference translation. Finally, the generated difficult words and interpretations can be assembled with the source and target sentence as demonstrations for each step of xIoD.
@@ 3.5 Demonstrations Synthesis for xIoD
| Zero-shot | 74.89 | 1.76 | 71.27 | -1.34 | 80.67 | 7.55 | 74.93 | 5.45 | 71.17 | -2.73 | 76.22 | 2.70 | 74.86 | 2.23 |
| ICL | 75.47 | 2.31 | 72.22 | -0.73 | 80.9 | 9.37 | 79.40 | 6.56 | 73.19 | -1.17 | 77.52 | 3.55 | 76.45 | 3.32 |
| +CoT | 73.85 | 0.86 | 71.35 | -1.00 | 78.03 | 5.24 | 76.78 | 4.77 | 69.72 | -2.96 | 76.55 | 2.96 | 74.38 | 1.65 |
| +Topic | 75.83 | 2.40 | 72.46 | -0.25 | 80.98 | 7.19 | 79.20 | 6.59 | 72.77 | -1.58 | 76.49 | 2.80 | 76.29 | 2.86 |
| +Keywords | 73.93 | 1.02 | 71.22 | -1.62 | 78.63 | 5.87 | 77.79 | 5.48 | 70.33 | -3.40 | 74.55 | 1.48 | 74.41 | 1.47 |
| +SimDems | 75.22 | 2.06 | 72.20 | -0.60 | 81.24 | 7.90 | 79.11 | 6.76 | 72.70 | -1.52 | 76.78 | 3.01 | 76.21 | 2.94 |
@@
| xIoD-I | 76.92 | 4.07 | 72.94 | 0.45 | 82.92 | 9.68 | 79.96 | 7.19 | 76.64 | 2.98 | 78.45 | 4.65 | 77.97 | 4.84 |
| xIoD-E | 77.57 | 4.61 | 73.23 | 0.32 | 83.07 | 10.39 | 80.01 | 7.81 | 77.04 | 3.03 | 78.70 | 4.82 | 78.27 | 5.16 |
4 Testbed: Challenge-MT dataset
In this work, we propose a benchmark Challenge-MT, which consists of difficult translation samples, for the following reasons: 1) MT has witnessed human-like performance on many languages, which appeals to a more challenging benchmark to detect the flaws of SOTA systems, and 2) a benchmark containing more instances that models underperform helps us to analyze the issue of the understanding misalignment in LLMs effectively.
Challenge-MT is constructed by collecting the most challenge subset of the widely used WMT datasets, involving six translation directions of Chinese (zh), Estonian (et), and Icelandic (is) to/from English (en). Specifically, we first evaluate three SOTA MT systems (Google Translate, ChatGPT (gpt-3.5-turbo), and NLLB NLLB Team et al. (2022)) on the multi-year WMT datasets. Due to the poor performance of NLLB on the zhen translation, we additionally train a zhen translation model based on the DeltaLM Ma et al. (2021) on the parallel corpus from OPUS††https://opus.nlpl.eu/. Next, The generated translations are scored with the COMET metric, and the of instances with the lowest score for each system are extracted as its difficult samples set. We vary the value of across different language pairs to ensure an appropriate scale for each difficult sample set. Finally, the intersection of all systems’ difficult sample sets is taken as the Challenge-MT benchmark. We equally split this dataset into the validation set and the test set.
As the results demonstrated in Fig 3, SOTA MT systems show extremely poor performance on the Challenge-MT benchmark, which is 10 COMET scores lower than the complete set in most translation directions. To understand the cause of the low performance on Challenge-MT, we conduct a multi-aspect comparison for the complete set and the Challenge-MT subset in Appendix A, which shows that the samples are longer and have higher perplexity in the Challenge-MT. These features indicate that sentences that are longer and harder to understand are more difficult to translate.
@@ 5.1 Experimental Setup
We verify the effectiveness of our xIoD on the LLM GPT-3.5-turbo for its promising capability in following complicated instructions. Demonstrations of xIoD are gained by performing our automatic method (§3.5) on the validation set of Challenge-MT. We compare xIoD with the following methods:
@@
For xIoD and other ICL-based methods, we select =8 demonstrations (i.e., 8-shot) to achieve a strong baseline performance. More details of re-implementing the baselines under the few-shot setting are illustrated in Appendix B
@@
Following previous research of LLM-based MT Garcia et al. (2023); Chen et al. (2023a), we adopt COMET Rei et al. (2020) as the evaluation metric as its higher correlation with human judgment than BLEU Papineni et al. (2002). Besides, we report the QE score to alleviate the reference bias Freitag et al. (2020) of the reference-based metrics.
5.2 Results

### v2

@@ Abstract
Large Language models (LLMs) have exhibited remarkable abilities in understanding complex texts, offering a promising path towards human-like translation performance. However, this study reveals the misalignment between the translation-specific understanding and the general understanding inside LLMs. This understanding misalignment leads to LLMs mistakenly or literally translating some complicated concepts that they accurately comprehend in the general scenarios (e.g., QA). To align the translation-specific understanding to the general one, we propose a novel translation process, DUAT (Difficult words Understanding Aligned Translation), explicitly incorporating the general understanding on the complicated content incurring inconsistent understanding to guide the translation. Specifically, DUAT performs cross-lingual interpretation for the difficult-to-translate words and enhances the translation with the generated interpretations. Furthermore, we reframe the external tools to improve DUAT in detecting difficult words and generating helpful interpretations. We conduct experiments on the self-constructed benchmark Challenge-WMT111The dataset is available at: ChallengeWMT 🖂 means corresponding author., consisting of samples that are prone to mistranslation. Human evaluation results on high-resource and low-resource language pairs indicate that DUAT significantly facilitates the understanding alignment, which improves the translation quality (up to +3.85 COMET) and reduces the literality of the translation by - 25% - 51%.
UTF8gbsn
@@
Yichong Huang†, Baohang Li†, Xiaocheng Feng†‡, Wenshuai Huo†‡, Chengpeng Fu†‡, Ting Liu†, Bing Qin †Harbin Institute of Technology ‡ Peng Cheng Laboratory {ychuang,xcfeng,baohangli,cpfu,wshuo,tliu,qinb}@ir.hit.edu.cn
@@ 1 Introduction
Recently, large language models (LLMs) have demonstrated remarkable language understanding and generation, paving the way for a higher level of performance in machine translation Zhao et al. (2023); OpenAI (2023); Jiang et al. (2023); Workshop (2023). However, existing research reports that LLMs have yet to achieve as significant advances in machine translation as they have achieved in other natural language processing fields Hendy et al. (2023); Pang et al. (2024); Zhang et al. (2023a); Jiao et al. (2023); Zhu et al. (2023b); Lou et al. (2023).
In this study, we discover the misalignment between the general understanding and translation-specific understanding inside LLMs, as illustrated in Fig.1. This understanding misalignment leads to LLMs mistakenly or literally translating some complicated concepts that they accurately comprehend in general scenarios. We refer to these failures as language models’ generalization failures on translation. Human evaluation on a total of 600 sampled sentences across six language pairs show that generalization failures account for a considerable proportion of all mistranslations (16%-32%), indicating serious understanding misalignment (§6.1).
@@
To align the translation-specific understanding to the general one, we propose a novel translation process, DUAT (Difficult words Understanding Aligned Translation), explicitly incorporating the general understanding on the complicated content incurring inconsistent understanding to guide the translation. Specifically, DUAT first detects the difficult-to-translate words in the source sentence, which could cover the generalization failures intuitively. Next, the LLM is prompted to interpret each difficult word with the target language, i.e., cross-lingual interpretation, unleashing the powerful general understanding and transforming this understanding into the target language space. After that, DUAT conducts translation under the guidance of these interpretations. Unlike the CoT-based process mimicking junior translators to sequentially translate all words Peng et al. (2023), DUAT works like senior translators to analyze the complicate words, which helps the model deep understand the source sentence and produces more nuanced translations. Furthermore, we reframe the external tool of token-level QE Rei et al. (2023) to enhance the detection of difficult words, and design a strategy of interpretation quality control to filter hallucinated interpretations based on sentence-level QE Rei et al. (2020).
To better analyze the understanding misalignment, we proposed the Challenge-WMT benchmark, which contains more sentences prone to mistranslation. These sentences were collected from multi-year WMT datasets and represent difficult samples that multiple state-of-the-art (SOTA) systems translate poorly. Human evaluation results indicate that DUAT significantly facilitates the understanding alignment, reducing 80%88% of generalization failures. Moreover, this alignment improves the translation quality, as evidenced by automatic metrics (up to +3.85 COMET), and alleviates translation literalness by -25% -51%.
@@ 2.1 LLM-based MT
2.2 Quality estimation (QE)
QE for machine translation, i.e., reference-free MT evaluation, aims to predict the quality of the given translation only according to the source sentence, which has shown auspicious correlations with human judgments Rei et al. (2020, 2021). Given a source sentence and a translation , QE score is denoted as .
@@ 2.2 Quality estimation (QE)
3 Approach: DUAT
In this section, we first introduce our translation framework DUAT (§3.1). Specifically, DUAT consists of three components: difficult word detection (§3.2), cross-lingual interpretation (§3.3), and interpretation quality control (§3.4). To make the LLM follow the procedure of each component as expected, we adopt the in-context learning strategy and design an automatic method for constructing demonstrations of DUAT (§3.5).
@@ 3.1 Framework
The progress of DUAT is illustrated in Fig.2. Given the source sentence, DUAT first detects the difficult words or phrases in the source sentence. Once the difficult words are identified, DUAT requests the LLM to interpret each difficult word with the target language, unleashing the powerful understanding capability inside the LLM and transforming these understandings into the target language space. Finally, to avoid the interference of incorrect and useless interpretations, DUAT removes the negative interpretations through the interpretation quality control and outputs the final translation guided by the helpful interpretations.
@@ 3.2 Step-1: Difficult Word Detection
In practice, we found that directly inquiring LLMs to identify difficult-to-translate words is challenging. To tackle this challenge, we first conduct a preliminary translation for the given source sentence and then extract the mistranslated words and phrases in the source sentence as the difficult words. Concretely, we invent DUAT-I to do this leveraging the Intrinsic ability of LLMs at first.
DUAT-I.
Given source sentence , DUAT-I first obtains the preliminary translation (also known as draft translation) by prompting the LLM to translate with the in-context learning strategy, which is shown in Eq. (1). Next, the LLM is requested to output the difficult words based on the source sentence and the preliminary translation:
DUAT-I obtains the difficult word list via performing greedy decoding on the LLM:
@@
DUAT-E.
It is frequently observed that the LLM fails to recognize the mistranslated words, due to their limitation in self-knowledge Yin et al. (2023) Therefore, we devise DUAT-E to boost the detection with the External tool. First, DUAT-E requests the LLM with the same prompt as DUAT-I while performing temperature sampling for times. Next, the union of all sampling results is taken as the candidate set of difficult words :
where is the hyperparameter of sampling temperature, which is set to to capture more candidates, and is set to 5 empirically.
Finally, DUAT-E annotates each candidate word with its degree of misalignment with respect to the draft translation, which reflects the translation-specific difficulty. To implement this function, we adopt an external tool of token-level QE . As shown in §2.2, token-level QE is originally used to score the mistranslation degree of the given translation span with respect to the source sentence, i.e., where . Differently, we utilize this tool in a dual manner. That is, we use to annotate the misalignment degree of the given source span with respect to the translation, i.e., where . Formally, the misalignment score of each difficult word candidate is calculated as:
Then, DUAT-E selects candidates with misalignment score , where is the hyperparameter named the difficulty threshold. We refer to this procedure as the difficulty-aware selection in Fig.2.
@@ 3.3 Step-2: Cross-Lingual Interpretation
After the difficult words in the source sentence are detected, DUAT lets the LLM generate the interpretation of each difficult word via requesting:
@@
Under the guidance of generated interpretations, DUAT can align the translation-specific understanding to the general one. However, LLMs may generate hallucinated interpretations sometimes (e.g., the interpretation of "崩塌" in Fig.2), which biases the resulting translation from the original semantics. Besides, helpless interpretations that can not provide useful information also pose a risk of disturbing the translation process.
@@ 3.4 Step-3: Interpretation Quality Control
To overcome the potential interference of the generated negative interpretations, DUAT removes them through the interpretation quality control (IQC) and outputs the final translation guided by the helpful interpretations.
Concretely, given a set of interpretations , DUAT ablates each interpretation sequentially and uses the remaining interpretations to guide the translation. The interpretation-guided translation is implemented in a fashion of refinement:
@@
where , which is the demonstration set of interpretation-guide translation.
If the better translation performance is achieved by ablation, which is measured by the QE222We use wmt21-comet-qe-da as the QE scorer. tool due to the unavailable access to the reference translation, the interpretation is removed from and the current translation is taken as the best translation. We also detail this process in Alg.1.
3.5 Demonstrations Synthesis for DUAT
To make the LLM follow the procedure of DUAT as expected, we adopt the ICL strategy. Common practice constructs demonstrations manually, necessitating human translators proficient in language pairs for languages. To overcome this considerable cost, we devise a method for synthesizing high-quality demonstrations of DUAT based on parallel data. This process of synthesizing demonstrations is accomplished in a manner of post-explanation by asking the LLM to compare the baseline translation and the reference translation. We describe this process in Appendix B.2.
@@ 3.5 Demonstrations Synthesis for DUAT
COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | COMET | BLEURT | |
@@
Zero-shot | 74.89 | 54.25 | 71.27 | 58.24 | 80.67 | 69.10 | 74.93 | 61.75 | 71.17 | 53.12 | 76.22 | 64.05 | 74.86 | 60.09 |
ICL | 75.47 | 55.79 | 72.22 | 59.56 | 80.9 | 69.93 | 79.40 | 66.63 | 73.19 | 54.87 | 77.52 | 65.81 | 76.45 | 62.10 |
+CoT | 73.85 | 53.42 | 71.35 | 57.90 | 78.03 | 66.03 | 76.78 | 63.97 | 69.72 | 50.98 | 76.55 | 64.54 | 74.38 | 59.47 |
+Keywords | 73.93 | 55.10 | 71.22 | 59.18 | 78.63 | 70.01 | 77.79 | 66.30 | 70.33 | 54.31 | 74.55 | 64.40 | 74.41 | 61.55 |
+Topic | 75.83 | 53.60 | 72.46 | 57.98 | 80.98 | 66.78 | 79.20 | 64.64 | 72.77 | 51.98 | 76.49 | 62.06 | 76.29 | 59.51 |
+SimDems | 75.22 | 55.10 | 72.20 | 59.16 | 81.24 | 70.42 | 79.11 | 66.29 | 72.70 | 54.04 | 76.78 | 64.43 | 76.21 | 61.57 |
@@
DUAT-I | 76.92 | 56.16 | 72.94 | 59.94 | 82.92 | 72.19 | 79.96 | 67.05 | 76.64 | 57.71 | 78.45 | 66.80 | 77.97 | 63.31 |
DUAT-E | 77.57 | 56.86 | 73.25 | 60.16 | 83.07 | 72.30 | 80.01 | 66.91 | 77.04 | 57.38 | 78.70 | 66.93 | 78.27 | 63.42 |
4 Testbed: Challenge-WMT
To better analyze the understanding misalignment problem of LLMs, we propose the benchmark Challenge-WMT, which contains challenge sentences that are prone to mistranslation. This benchmark is constructed by collecting samples that multiple SOTA systems translate poorly from multi-year WMT testsets of six language pairs (Chinese (zh), Estonian (et), and Icelandic (is) to/from English (en)). Additionally, we believe that this dataset could promote future research in understanding the limitations of existing MT systems.
We select three SOTA MT systems: Google Translate, ChatGPT, and NLLB NLLB Team et al. (2024). Due to the poor performance of NLLB in the zhen translation, we additionally train a zhen translation model based on DeltaLM Ma et al. (2021) on the parallel corpus from OPUS333https://opus.nlpl.eu/. Next, all of the system translations are scored with COMET metric, and the of samples with the lowest score for each system are extracted as its difficult samples set. We vary the value of across different language pairs to ensure an appropriate scale for each difficult sample set. Finally, the intersection of all systems’ difficult sample sets is taken as the Challenge-WMT testbed. Challenge-WMTcomprises around 600+ sentence pairs for each language pair, which is illustrated in Tab.6. We equally split this dataset into the validation set and the test set.
We report the translation performance measured by COMET on Challenge-WMT and the complete WMT set in Fig.8, which shows that the performance decreases dramatically on Challenge-WMT (84.573.6 averagely). Next, we conduct a multi-aspect comparison for Challenge-WMT and the complete set in Appendix C, and find that the samples of Challenge-WMT have higher perplexity (214252 averagely). This result indicates that sentences in Challenge-WMT are more complex.
@@ 5.1 Experimental Setup
We verify the effectiveness of our DUAT on the LLM GPT-3.5-turbo for its promising capability in following complicated instructions. Demonstrations of DUAT are gained by performing our automatic method (§3.5) on the validation set of Challenge-WMT. We compare DUAT with the following methods:
@@
For DUAT and other ICL-based methods, we select =8 demonstrations (i.e., 8-shot) to achieve a strong baseline performance. More details of re-implementing the baselines under the few-shot setting are illustrated in Appendix D.
@@
5.2 Results on Challenge-WMT